# Magnetocalorics of singlet ground state induced moment magnets

### v1

@@ V The case of PrI3: a near-critical singlet induced ferromagnet
As noted in the introduction the singlet-singlet CEF models are frequently realized in compounds of 4f-Pr- and 5f-U- transition elements in non-Kramers configuration ($f^{2}$, $J=4$) in a suitable low- symmetry CEF environment. A recent example is the trihalide compound PrI3 described in Ref. 16. It exhibits low temperature ferromagnetic order at $T_{m}\equiv T_{c}=7.5\;\text{K}$. Due to the low site symmetry $C_{2}$ of Pr ions in the low temperature phase only singlet CEF states can be present. Then the FM ordering should be of the induced moment type discussed here and previously Ref.15. Indeed the observed anomaly of the specific heat (sofar only in zero field) around the ordering temperature supports such interpretation. It shows the two signatures of the induced moment magnet as evident from the experimental data in Fig. 6 adapted from Ref. 16: Firstly an underlying Schottky type anomaly resulting from the single-site CEF singlet-singlet splitting. Secondly the specific heat jump associated with the induced ferromagnetic order. As explained before the relative position of the maximum $(T^{C_{V}}_{max})$ and the ordering temperature $T_{m}$ depends on the control parameter $\xi_{s}$. In the case of PrI3 it was observed that the maximum $(T^{\gamma_{V}}_{max})$ of $\gamma_{V}=C_{V}(T)/T$ coincides with the ordering temperature. Then, following the arguments in the previous section this leads to values $\Delta=24.4\;\text{K}$ for the singlet splitting and a control parameter $\xi_{s}=1.08$ for PrI3. Therefore the new compound PrI3 may be viewed as a induced singlet ground state ferromagnet rather close to the QCP at $\xi_{s}=1$. In fact when plotting $C(T)$ itself for $\xi_{s}=1.08$ the specific heat jump at $T_{m}=7.5\;\text{K}$ lies already considerably below the Schottky maximum.
@@
For this asymptotic case with $\xi_{s}=1+\delta$ ($\delta\ll 1$) the relative jump size vanishes like [15] $(C_{V}^{-}-C_{V}^{+})/C_{V}^{+}\approx\delta|ln\frac{\delta}{2}|$ where $C_{V}^{\pm}$ are the values close above and below the transition. We finally note that the case where specific heat jump at $T_{m}$ is directly on top of the Schottky anomaly at $T_{max}$ of $C_{V}(T)$ itself occurs for $\xi_{s}\simeq 1.2$. For control parameters around this value entropy release by ordering and level depopulation may be considered as comparatively important (Fig. 8). This value of $\xi_{s}$ has mostly been adopted in the figures which are not directly related to PrI3.
@@ VI Field dependence of magnetic exciton modes
In the molecular field picture the inter-site exchange in the singlet-singlet system leads to possible induced order with spontaneous moment creation due to singlet mixing and a concomitant renormalization of the CEF splitting of singlets, both depending on the external field as well. Considering the effect of the exchange beyond this static picture, i.e., in an RPA response function approach [3, 15] allows to study the dynamics of the model. The finite CEF singlet gap then turns into dispersive magnetic exciton modes due to the inter-site exchange which are already present in the disordered phase. They have been investigated experimentally [22, 4, 23, 6, 7] as well as theoretically [24, 5, 25, 15] for numerous singlet ground state magnets. Their temperature dependence is characterized by (mostly arrested) soft mode behavior above the ordering temperature around the wave vector of the incipient induced magnetic order. As shown in detail in Ref. [15] the latter depends crucially on the nature of the first excited state, whether it is a singlet, doublet or triplet.
@@
$\chi_{0\lambda}(i\omega_{n},T,h^{\prime})$ $=$ $m_{s}^{2}\sum_{\tau\neq\tau^{\prime}}|\langle\tau|S_{x}|\tau^{\prime}\rangle|^{2}_{\lambda}\frac{p^{\tau}_{\lambda}-p^{\tau^{\prime}}_{\lambda}}{\epsilon^{\tau^{\prime}}_{\lambda}-\epsilon^{\tau}_{\lambda}-i\omega_{n}}$ (43)
where $\tau=\pm$ denote the mixed singlet eigenstates on sublattice $\lambda=A,B$. For finite frequency this contains only the vanVleck term without the zero-frequency pseudo-Curie term in the static case (Eq.17). It may be evaluated as
$\chi_{0\lambda}(i\omega_{n})=\bigl{(}\frac{m_{s}^{2}}{2\hat{\Delta}^{\lambda}_{T}}\bigr{)}\frac{\Delta}{(\Delta\hat{\Delta}_{T}^{\lambda})^{2}-(i\omega_{n})^{2}}f_{s}^{\lambda}(T)$ (44)
@@
$\omega_{F}({\bf q},T,h^{\prime})=\Delta\hat{\Delta}_{T}\bigl{[}1-\frac{\xi_{s}}{\hat{\Delta}_{T}^{3}}f_{s}(T)\gamma({\bf q})\bigr{]}^{\frac{1}{2}}$ (45)
@@
$\omega_{0}({\bf q},T,h^{\prime})=\Delta\hat{\Delta}_{T}\bigl{[}1+\frac{\xi_{s}}{\hat{\Delta}_{T}^{3}}f_{s}(T)\gamma({\bf q})\bigr{]}^{\frac{1}{2}}$ (46)
@@
±$\pm\bigl{[}\frac{1}{4}(\hat{\Delta}_{T}^{A2}-\hat{\Delta}_{T}^{B2})^{2}+\xi_{s}^{2}\frac{f_{s}^{A}(T)f_{s}^{B}(T)}{\hat{\Delta}_{T}^{A}\hat{\Delta}_{T}^{B}}\gamma({\bf q})^{2}\bigr{]}^{\frac{1}{2}}\Bigr{\}}^{\frac{1}{2}}$
@@
In the FM case Eq. (45) is appropriate in the whole temperature and field range with $\hat{\Delta}_{T}$ obtained from Eq. (10) and $f_{s}(T)$ from Eq. (15). For zero field a true phase transition occurs for $\xi_{s}>1$. For $T\rightarrow T_{m}$ we have $\Delta_{T}\rightarrow 1$ and then $f_{s}(T)\rightarrow\xi_{s}^{-1}$. Furthermore at the FM ordering vector ${\bf q}=0$ we get $\gamma(0)=1$. Then Eq. (45) leads to $\omega(0,T_{m},0)\rightarrow 0$, i.e. a soft mode at the FM $\Gamma$-point appears as a precursor to the induced FM ordering. For finite field when only a crossover to a FM polarized state occurs the soft mode acquires a finite frequency again already at small field sizes.
The soft mode precursor behaves even more interesting in the AFM phase transition. For $T<T_{m}(h^{\prime})$ there are now generally two modes from Eq. (VI) because $\hat{\Delta}_{T}^{A}\neq\hat{\Delta}_{T}^{B}$ on the two sublattices. In the paramagnetic phase with $\hat{\Delta}_{T}^{\lambda}=\hat{\Delta}_{T}$ it reduces to the same form as Eqs. (45,VI) but with $\omega_{\pm}({\bf q})$ corresponding to a downfolding from the simple cubic paramagnetic Brillouin zone (BZ) to the fcc antiferromagnetic BZ (AFBZ). The $1$ AFBZ is defined by the boundaries ${\bf q}\cdot{\bf Q}^{\prime}=\pm 1$ where ${\bf Q}^{\prime}$ is any of the eight midpoint (L) vectors (±$(\pm\frac{1}{2},\pm\frac{1}{2},\pm\frac{1}{2})$. The unfolding in the paramagnetic phase is defined by $\omega_{+}({\bf q})=\omega_{F}({\bf q})$ and $\omega_{-}({\bf q})=\omega_{0}({\bf q})$ in the $2$ AFBZ and vice versa in the $1$ AFBZ (the complement in the simple cubic BZ) as further discussed in Sec. VII.
With similar arguments as above for $\xi_{s}>1$ and zero field the mode frequencies at the AFM ordering vector ${\bf Q}=(1,1,1)$ (R-point) are $\omega_{\pm}({\bf Q},T_{m},0)=0,2\Delta$ with $\omega_{-}({\bf Q})$ now corresponding to the precursor soft mode. In the AFM case the transition at $T_{m}(h^{\prime})$ exists also for finite fields $h^{\prime}<h^{\prime}_{c}$ below the critical one. Then the question arises whether the zero-field soft mode appearance at $T_{m}$ seen before persists also for finite field for all $T_{m}(h^{\prime})$. This is, however, not the case for the following reason: The AFM critical temperature as determined by Eq. (29) contains also the effect of the (second) zero-frequency pseudo-Curie terms leading to a slightly larger $T_{m}(h)$ than to be expected from the (first) vanVleck terms only, However the mode frequency is determined solely by the poles due to the dynamical vanVleck terms and therefore the AFM transition $T_{m}(h^{\prime})$ happens already before the complete softening of the precursor mode takes place. This ’arrested soft mode’ frequency may be obtained from Eq. (VI) as
@@
$\Delta\bigl{(}\frac{\xi_{s}\Delta}{2T_{m}}\bigr{)}^{\frac{1}{2}}\bigl{(}\hat{\Delta}^{2}_{T_{m}}-1\bigr{)}^{\frac{1}{2}}(1-f_{s}^{2}(T_{m}))^{\frac{1}{2}}\geq 0$ (48)
which vanishes in the zero field limit where $\hat{\Delta}_{T_{m}}=1$ and also at the critical field where $T_{m}(h^{\prime}_{c})\rightarrow 0$ leading to $f_{s}(T_{m})\rightarrow 1$. Therefore strictly, even in a singlet-singlet model a true soft mode occurs only at $T_{m}(h^{\prime})$ for $h^{\prime}=0,h_{c}^{\prime}$ while an ’arrested’ soft mode with $\omega_{-}({\bf Q},T_{m}(h^{\prime}),h^{\prime})>0$ is observed in the interval $0<h^{\prime}<h^{\prime}_{c}$. (We note that e.g. in a singlet-triplet induced moment magnet it is already present at zero field due to the effect of the genuine Curie terms in the static susceptibility resulting from the excited triplet [26, 15]). This naively unexpected behavior is clearly seen in Figs. 2(a) and Fig. 12.
Figure 11: Magnetic exciton dispersions for $\xi_{s}=1.2$ in the simple cubic BZ along $X(010),\Gamma(000),R(111),M(110),\Gamma(000)$ path. This is an extended zone scheme with respect to the two-sublattice AFM (d) ordered phase. (a) FM zero-field critical case at the transition temperature $T_{m}/\Delta=0.417$ with soft mode at $\Gamma$ point. (b) FM again at $T_{m}$ for finite field $h/\Delta=0.05$ with rapid gapping of the soft mode at $\Gamma$. (c) AFM zero-field critical case at the transition temperature $T_{m}/\Delta=0.417$ with soft mode at R(111) point. (d) AFM at finite field ($h/\Delta$=0.6 and $T/\Delta$ =0.3 ) inside the ordered phase. The soft mode will be gapped and hybridization of upper and lower modes occur (in the extended zone scheme of the AFM R and $\Gamma$ are equivalent and the L(12,12,$(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ midpoint of $\Gamma$-R segment is an AFBZ zone boundary point). For soft mode field dependence along $T_{m}(h)$ transition line see Fig. 2. A logarithmic scale for better visibility of dispersions has been employed for the mode intensity.
@@
$S({\bf q},\omega)=\frac{1}{\pi}Im\sum_{\lambda}\hat{\chi}_{\lambda\lambda}({\bf q},i\omega_{n})_{i\omega_{n}\rightarrow\omega+i\eta}$ (49)
@@
$FM:\;\;S({\bf q},\omega)=\frac{1}{2}m_{s}^{2}\frac{\hat{f}_{s}(T)}{2\hat{\omega}_{\bf q}}\delta(\omega-\omega_{\bf q})$ (50)
@@
$\hat{f}_{sn}(T)$ $=$ $\frac{\sum_{\lambda}\hat{f}_{s}^{\lambda}(T)(\hat{\omega}^{2}_{n{\bf q}}-\hat{\Delta}^{\bar{\lambda}2}_{T})}{(\hat{\omega}^{n2}_{\bf q}-\hat{\omega}^{\bar{n}2}_{\bf q})}$ (51)
where $n=\pm$; $(\bar{n}=\mp)$ denotes the two modes and $\lambda=A,B$; $(\bar{\lambda}=B,A)$ is as before the sublattice index. Furthermore as above we introduced $\hat{f}_{s}^{\lambda}(T)=f_{s}^{\lambda}(T)/\hat{\Delta}^{\lambda}_{T}$ (cf. Eq.(13)) and $\hat{\omega}_{n{\bf q}}=\omega_{\bf q}/\Delta$. It is obvious from the FM case but also true for AFM that the mode intensity increases with decreasing frequency and for the soft mode diverges (Fig.12). The mode dispersions and spectral functions for some typical and instructive choices of field and temperature in ordered and paramagnetic phases are shown in Figs. 11,12 and their implications are explained below.

### v2

@@ V.1 The case of PrIr3 : a new near-critical singlet induced ferromagnet
Now we consider a most recent addition to these examples in more detail, in the context of the previous theoretical analysis, which is the intermetallic compound PrIr3 described in Ref. 4. It exhibits low temperature ferromagnetic order at $T_{m}\equiv T_{C}=7.5\;\text{K}$. Because of the low site symmetry $C_{2}$ of Pr ions in the low temperature phase only singlet CEF states can be present. Due to its intermetallic nature the effective intersite interaction between the Pr3+ localised 4f moments is dominated by the RKKY mechanism mediated by conduction electrons [10]. Then the FM ordering should be of the induced moment type discussed here and previously [3]. Indeed the observed anomaly of the specific heat (sofar only in zero field) around the ordering temperature supports such interpretation. It shows the two signatures of the induced moment magnet as evident from the experimental data in Fig. 5 adapted from Ref. 4: Firstly an underlying Schottky type anomaly resulting from the single-site CEF singlet-singlet splitting. Secondly the specific heat jump associated with the induced ferromagnetic order. As explained before the relative position of the maximum $(T^{C_{V}}_{max})$ and the ordering temperature $T_{m}$ depends on the control parameter $\xi_{s}$. In the case of PrIr3 it was observed that the maximum $(T^{\gamma_{V}}_{max})$ of $\gamma_{V}=C_{V}(T)/T$ coincides with the ordering temperature. Then, following the arguments in the previous section this leads to values $\Delta=24.4\;\text{K}$ for the singlet splitting and a control parameter $\xi_{s}=1.08$ for PrIr3. Therefore the new compound PrIr3 may be viewed as a induced singlet ground state ferromagnet rather close to the QCP at $\xi_{s}=1$. In fact when plotting $C(T)$ itself for $\xi_{s}=1.08$ the specific heat jump at $T_{m}=7.5\;\text{K}$ lies already considerably below the Schottky maximum. For this asymptotic case with $\xi_{s}=1+\delta$ ($\delta\ll 1$) the relative jump size vanishes like [3] $(C_{V}^{-}-C_{V}^{+})/C_{V}^{+}\approx\delta|ln\frac{\delta}{2}|$ where $C_{V}^{\pm}$ are the values close above and below the transition [3].
We stress that we do not intend a quantitative comparison of specific heat temperature dependences because PrIr3  has a linear low temperature itinerant specific heat contribution (subtracted in Fig. 5) and also presumably contributions from higher CEF states at elevated temperatures (the entropy exceeds $ln2$ there), both effects are not described by our two-singlet model. But the comparison in Fig. 5 demonstrates a similarity of shapes for $C_{V}/T$: A jump due to FM order sitting right on top of the Schottky peak. The essential point we want to make is that the relative position $T_{max}$ of low temperature Schottky background maximum and ordering temperature Tm gives information on the control parameter and the associated distance to the FM QCP of the induced moment magnetism (see also Fig. 6).
@@
We finally note that the case where specific heat jump at $T_{m}$ is directly on top of the Schottky anomaly at $T_{max}$ of $C_{V}(T)$ itself occurs for $\xi_{s}\simeq 1.2$. For control parameters around this value entropy release by ordering and level depopulation may be considered as comparatively important (Fig. 7). This value of $\xi_{s}$ has mostly been adopted in the figures which are not directly related to PrIr3.
@@ VI Field dependence of magnetic exciton modes
In the molecular field picture the inter-site exchange in the singlet-singlet system leads to possible induced order with spontaneous moment creation due to singlet mixing and a concomitant renormalisation of the CEF splitting of singlets, both depending on the external field as well. Considering the effect of the exchange beyond this static picture, i.e., in an RPA response function technique [10, 3] allows to study the dynamics of the model. An alternative for low temperatures is provided by using the Bogoliubov transformation approach [3]. The finite CEF singlet gap then turns into dispersive magnetic exciton modes due to the inter-site exchange which are already present in the disordered phase. They have been investigated experimentally [19, 18, 20, 21, 22] as well as theoretically [46, 26, 7, 3] for numerous singlet ground state magnets. Their temperature dependence is characterized by (mostly arrested) soft mode behavior above the ordering temperature around the wave vector of the incipient induced magnetic order. As shown in detail in Ref. [3] the mode softening depends crucially on the nature of the first excited state, whether it is a singlet, doublet or triplet.
@@
$\hskip-14.22636pt\chi_{0\lambda}(i\omega_{n},T,h^{\prime})$ $=$ $m_{s}^{2}\sum_{\tau\neq\tau^{\prime}}|\langle\tau|S_{x}|\tau^{\prime}\rangle|^{2}_{\lambda}\frac{p^{\tau}_{\lambda}-p^{\tau^{\prime}}_{\lambda}}{\epsilon^{\tau^{\prime}}_{\lambda}-\epsilon^{\tau}_{\lambda}-i\omega_{n}},$ (43)
where $\tau=\pm$ denote the mixed singlet eigenstates on sublattice $\lambda=A,B$. For finite frequency this contains only the vanVleck term without the zero-frequency pseudo-Curie term in the static case (Eq.(17)). It may be evaluated as
$\chi_{0\lambda}(i\omega_{n})=\bigl{(}\frac{m_{s}^{2}}{2\hat{\Delta}^{\lambda}_{T}}\bigr{)}\frac{\Delta}{(\Delta\hat{\Delta}_{T}^{\lambda})^{2}-(i\omega_{n})^{2}}f_{s}^{\lambda}(T).$ (44)
@@
$\omega_{F}({\bf q},T,h^{\prime})=\Delta\hat{\Delta}_{T}\bigl{[}1-\frac{\xi_{s}}{\hat{\Delta}_{T}^{3}}f_{s}(T)\gamma({\bf q})\bigr{]}^{\frac{1}{2}},$ (45)
@@
$\omega_{0}({\bf q},T,h^{\prime})=\Delta\hat{\Delta}_{T}\bigl{[}1+\frac{\xi_{s}}{\hat{\Delta}_{T}^{3}}f_{s}(T)\gamma({\bf q})\bigr{]}^{\frac{1}{2}},$ (46)
@@
$\pm\bigl{[}\frac{1}{4}(\hat{\Delta}_{T}^{A2}-\hat{\Delta}_{T}^{B2})^{2}+\xi_{s}^{2}\frac{f_{s}^{A}(T)f_{s}^{B}(T)}{\hat{\Delta}_{T}^{A}\hat{\Delta}_{T}^{B}}\gamma({\bf q})^{2}\bigr{]}^{\frac{1}{2}}\Bigr{\}}^{\frac{1}{2}}.$
@@
In the FM case Eq. (45) is appropriate in the whole temperature and field range with $\hat{\Delta}_{T}$ obtained from Eq. (10) and $f_{s}(T)$ from Eq. (15). For zero field a true phase transition occurs for $\xi_{s}>1$. For $T\rightarrow T_{m}$ we have $\hat{\Delta}_{T}\rightarrow 1$ and then $f_{s}(T)\rightarrow\xi_{s}^{-1}$. Furthermore at the FM ordering vector ${\bf q}=0$ we get $\gamma(0)=1$. Then Eq. (45) leads to $\omega(0,T_{m},0)\rightarrow 0$, i.e. a soft mode at the FM $\Gamma$-point appears as a precursor to the induced FM ordering. For finite field when only a crossover to a FM polarized state occurs the soft mode acquires a finite frequency again already at small field sizes.
The soft mode precursor behaves even more interesting in the AFM phase transition. For $T<T_{m}(h^{\prime})$ there are now generally two modes from Eq. (VI) because $\hat{\Delta}_{T}^{A}\neq\hat{\Delta}_{T}^{B}$ on the two sublattices. In the paramagnetic phase with $\hat{\Delta}_{T}^{\lambda}=\hat{\Delta}_{T}$ it reduces to the same form as Eqs. (45,VI) but with $\omega_{\pm}({\bf q})$ corresponding to a downfolding from the simple cubic paramagnetic Brillouin zone (BZ) to the fcc antiferromagnetic BZ (AFBZ). The $1.$ AFBZ is defined by the boundaries ${\bf q}\cdot{\bf Q}^{\prime}=\pm 1$ where $\bf Q$$^{\prime}$ is any of the eight midpoint (L) vectors $(\pm\frac{1}{2},\pm\frac{1}{2},\pm\frac{1}{2})$. The unfolding in the paramagnetic phase is defined by $\omega_{+}({\bf q})=\omega_{F}({\bf q})$ and $\omega_{-}({\bf q})=\omega_{0}({\bf q})$ in the $2.$ AFBZ and vice versa in the $1.$ AFBZ (the complement in the simple cubic BZ) as further discussed in Sec. VII.
With similar arguments as above for $\xi_{s}>1$ and zero field the mode frequencies at the AFM ordering vector ${\bf Q}=(1,1,1)$ (R-point) are $\omega_{\pm}({\bf Q},T_{m},0)=0,2\Delta$ with $\omega_{-}({\bf Q})$ now corresponding to the precursor soft mode. In the AFM case the transition at $T_{m}(h^{\prime})$ exists also for finite fields $h^{\prime}<h^{\prime}_{c}$ below the critical one. Then the question arises whether the zero-field soft mode appearance at $T_{m}$ seen before persists also for finite field for all $T_{m}(h^{\prime})$. This is, however, not the case for the following reason: The AFM critical temperature as determined by Eq. (29) contains also the effect of the (second) zero-frequency pseudo-Curie terms leading to a slightly larger $T_{m}(h)$ than to be expected from the (first) vanVleck terms only, However the mode frequency is determined solely by the poles due to the dynamical vanVleck terms and therefore the AFM transition $T_{m}(h^{\prime})$ happens already before the complete softening of the precursor mode takes place. This ’arrested soft mode’ frequency is obtained from Eq. (VI) as
@@
$\Delta\bigl{(}\frac{\xi_{s}\Delta}{2T_{m}}\bigr{)}^{\frac{1}{2}}\bigl{(}\hat{\Delta}^{2}_{T_{m}}-1\bigr{)}^{\frac{1}{2}}(1-f_{s}^{2}(T_{m}))^{\frac{1}{2}}\geq 0,$ (48)
which vanishes in the zero field limit where $\hat{\Delta}_{T_{m}}=1$ and also at the critical field where $T_{m}(h^{\prime}_{c})\rightarrow 0$ leading to $f_{s}(T_{m})\rightarrow 1$. Therefore strictly, even in a singlet-singlet model a true soft mode occurs only at $T_{m}(h^{\prime})$ for $h^{\prime}=0,h_{c}^{\prime}$ while an ’arrested’ soft mode with $\omega_{-}({\bf Q},T_{m}(h^{\prime}),h^{\prime})>0$ is observed in the interval $0<h^{\prime}<h^{\prime}_{c}$. (In a singlet-triplet induced moment magnet this happens already at zero field due to the effect of the genuine Curie terms in the static susceptibility resulting from the excited triplet [47, 3]). The naively unexpected behavior of $\omega_{-}({\bf Q},T_{m}(h^{\prime}),h^{\prime})$ is clearly seen in Figs. 2(a) and Fig. 11.
Figure 10: Magnetic exciton dispersions for $\xi_{s}=1.2$ in the simple cubic BZ along $X(010),\Gamma(000),R(111),M(110),\Gamma(000)$ path. This is an extended zone scheme with respect to the two-sublattice AFM (d) ordered phase. (a) FM zero-field critical case at the transition temperature $T_{m}/\Delta=0.417$ with soft mode at $\Gamma$ point. (b) FM again at $T_{m}$ for finite field $h/\Delta=0.05$ with rapid gapping of the soft mode at $\Gamma$. (c) AFM zero-field critical case at the transition temperature $T_{m}/\Delta=0.417$ with soft mode at R(111) point. (d) AFM at finite field ($h/\Delta$=0.6 and $T/\Delta$ =0.3 ) inside the ordered phase. The soft mode will be gapped and hybridization of upper and lower modes occur (in the extended zone scheme of the AFM R and $\Gamma$ are equivalent and the L(12,12,12)($\frac$$\frac{1}{2}$$\frac${1}{2}) midpoint of $\Gamma$-R segment is an AFBZ zone boundary point). For soft mode field dependence along $T_{m}(h)$ transition line see Fig. 2. A logarithmic scale for better visibility of dispersions has been employed for the mode intensity.
@@
$S({\bf q},\omega)=\frac{1}{\pi}Im\sum_{\lambda}\hat{\chi}_{\lambda\lambda}({\bf q},i\omega_{n})_{i\omega_{n}\rightarrow\omega+i\eta}.$ (49)
@@
$FM:\;\;S({\bf q},\omega)=\frac{1}{2}m_{s}^{2}\frac{\hat{f}_{s}(T)}{2\hat{\omega}_{\bf q}}\delta(\omega-\omega_{\bf q}),$ (50)
@@
$\hat{f}_{sn}(T)$ $=$ $\frac{\sum_{\lambda}\hat{f}_{s}^{\lambda}(T)(\hat{\omega}^{2}_{n{\bf q}}-\hat{\Delta}^{\bar{\lambda}2}_{T})}{(\hat{\omega}^{n2}_{\bf q}-\hat{\omega}^{\bar{n}2}_{\bf q})},$ (51)
where $n=\pm$; $(\bar{n}=\mp)$ denotes the two modes and $\lambda=A,B$; $(\bar{\lambda}=B,A)$ is as before the sublattice index. Furthermore as above we introduced $\hat{f}_{s}^{\lambda}(T)=f_{s}^{\lambda}(T)/\hat{\Delta}^{\lambda}_{T}$ (cf. Eq.(13)) and $\hat{\omega}_{n{\bf q}}=\omega_{\bf q}/\Delta$. It is obvious from the FM case but also true for AFM that the mode intensity increases with decreasing frequency and for the soft mode diverges (Fig.11). The mode dispersions and spectral functions for some typical and instructive choices of field and temperature in ordered and paramagnetic phases and their implications are explained below.